OpenStudy (anonymous):

A parachute can travel for 4 consecutive hours continuously using gas only. A glider can glide 14.5% longer using the same the thermals. If a glider makes 3 stops during a certain glide, each after gliding its maximum possible time, how many hours long was the glide not including the resting time? 18.32,13.74,15.66,9.16

OpenStudy (anonymous):

14.5% of 4 hours = 2088 seconds So the glider can glide for 16488 seconds per flight 16488 * 3 = 49464 seconds 49464 / 3600 = 13,74 hours

OpenStudy (anonymous):

parachute= 4 consecutive hours glider= 4+ (14.5%)(4) given: the exception of the resting time meaning: resting time is not included, therefor: find it's maximum time. . that's the process i can only think. . but you could correct that if it's wrong! lol

OpenStudy (anonymous):

ooh, i forgot! multiply by 3 the glider, since resting time is excepted.

OpenStudy (anonymous):

that's 13.74, if my calculator was not wrong! lol

OpenStudy (anonymous):

thnks to both of you

OpenStudy (anonymous):

medalsss. . .